OpenStudy (anonymous):

3x + 1 + 5x = 7 +15 + 7x solve for x .-.

OpenStudy (xmissalycatx):

Hello there! :)

OpenStudy (ineedhelplz):

Let's solve your equation step-by-step. 3x+1+5x=7+15+7x Step 1: Simplify both sides of the equation. 3x+1+5x=7+15+7x Simplify: (Show steps) 8x+1=7x+22 Step 2: Subtract 7x from both sides. 8x+1−7x=7x+22−7x x+1=22 Step 3: Subtract 1 from both sides. x+1−1=22−1 x=21 Answer: x=21

OpenStudy (anonymous):

lol thanks

OpenStudy (xmissalycatx):

Really? You're going to copy and paste it? AND you give a direct answer? SMH.

OpenStudy (anonymous):

i know what he did @xMissAlyCatx